OpenStudy (anonymous):

Find the zeroes of the function by factoring: f(x)= (64x^2-4)/x

OpenStudy (anonymous):

(8x+2)(8x-2) ------------ x

OpenStudy (anonymous):

to make it zero we have to see for what x value the following is zero (8x-2) x=1/4

OpenStudy (anonymous):

thus zero is x=1/4

OpenStudy (anonymous):

my confusion with this question is that since x cannot equal 0 in this equation, would that be a zero? im blanking haha sorry. and wouldnt the other zero also be x= -1/4?

OpenStudy (anonymous):

they meant for what value of x is the function is zero

OpenStudy (anonymous):

F(x)=x-2 for example , zero of above function is 2 because that makes F(x)=0

OpenStudy (anonymous):

OHHH okay right. thank you so much! summer work makes me feel stupid-_____-